Aspire's Library

A Place for Latest Exam wise Questions, Videos, Previous Year Papers,
Study Stuff for MCA Examinations

MCA NIMCET Previous Year Questions (PYQs)

MCA NIMCET Probability Distribution PYQ


MCA NIMCET PYQ
Probability a blade is defective $=0.002$, packet of $10$ blades. Find packets with no defective blades in $10000$ packets.





Go to Discussion

MCA NIMCET Previous Year PYQ MCA NIMCET NIMCET 2011 PYQ

Solution

$P(\text{good})=0.998$ Probability all 10 good $=(0.998)^{10}\approx 0.9802$ So packets $=10000\times0.9802\approx9802$

MCA NIMCET PYQ
An anti-aircraft gun fires a maximum of four shots. Probabilities of hitting in the 1st, 2nd, 3rd, and 4th shot are 0.4, 0.3, 0.2 and 0.1 respectively. Find the probability that the gun hits the plane.





Go to Discussion

MCA NIMCET Previous Year PYQ MCA NIMCET NIMCET 2009 PYQ

Solution

Hit at least once = 1 − (miss all shots) 
Miss probabilities: 
1st: $(1 - 0.4) = 0.6$ 
2nd: $(1 - 0.3) = 0.7$ 
3rd: $(1 - 0.2) = 0.8$ 
4th: $(1 - 0.1) = 0.9$ 
Miss all: $0.6 \cdot 0.7 \cdot 0.8 \cdot 0.9 = 0.3024$ 
Hit at least once: $1 - 0.3024 = 0.6976$


MCA NIMCET PYQ
A random variable $X$ has the probability distribution: \[\begin{array}{c|ccccccccc} x & 0 & 1 & 2 & 3 & 4 & 5 & 6 & 7 & 8 \\ \hline P(X=x) & a & 3a & 5a & 7a & 9a & 11a & 13a & 15a & 17a \end{array} \]The value of $a$ is:





Go to Discussion

MCA NIMCET Previous Year PYQ MCA NIMCET NIMCET 2011 PYQ

Solution

$ \text{Total probability} = 1 $ $ a(1 + 3 + 5 + 7 + 9 + 11 + 13 + 15 + 17) = 1 $ $ 1 + 3 + 5 + 7 + 9 + 11 + 13 + 15 + 17 = 81 $ $ 81a = 1 $ $ a = \frac{1}{81} $



MCA NIMCET PYQ
A random variable X has the distribution law as given below:
 X 123 
 P(X=x) 0.30.4 0.3 
The variance of the distribution is





Go to Discussion

MCA NIMCET Previous Year PYQ MCA NIMCET NIMCET 2013 PYQ

Solution


MCA NIMCET PYQ
India plays 4 matches. Probabilities of scoring $0,1,2$ points: $0.45,\ 0.05,\ 0.50$. Find probability of at least $7$ points.





Go to Discussion

MCA NIMCET Previous Year PYQ MCA NIMCET NIMCET 2010 PYQ

Solution

$P(8)=0.5^4=0.0625$ 
$P(7)=4 \cdot 0.05 \cdot 0.5^3 = 0.025$ 
Total $= 0.0625 + 0.025 = 0.0875$

MCA NIMCET PYQ
n a Poisson distribution if $P[X=3] = \frac{1}{4} P[X=4]$ then $P[X=5] = k P[X=7]$ where $k$ equals to





Go to Discussion

MCA NIMCET Previous Year PYQ MCA NIMCET NIMCET 2010 PYQ

Solution

Poisson PMF: 
$P[X=x] = e^{-\lambda} \frac{\lambda^x}{x!}$ 
Given: $P(3) = \frac{1}{4} P(4)$ 
 $\frac{\lambda^3/3!}{\lambda^4/4!} = \frac{1}{4}$ 
 $\frac{4}{\lambda} = \frac{1}{4}$ 
 $\lambda = 16$ 
 Now find $k$: 
$\frac{P(5)}{P(7)} = \frac{\lambda^5/5!}{\lambda^7/7!} = \frac{7 \cdot 6}{\lambda^2}$ 
 Substitute $\lambda = 16$: 
$\frac{42}{256} = \frac{21}{128}$ 
 So, $k = \frac{21}{128}$

MCA NIMCET PYQ
The mean and variance of a random variable X having binomial distribution are 4 and 2 respectively. The P(X = 1) is





Go to Discussion

MCA NIMCET Previous Year PYQ MCA NIMCET NIMCET 2017 PYQ

Solution

np = 4

npq = 2

q = 1/2, p = 1/2, n = 8

p(X = 1) = 8C1 (1/2)(1/2)7


MCA NIMCET PYQ
If  is the mean of distribution of x, then usual notation  is





Go to Discussion

MCA NIMCET Previous Year PYQ MCA NIMCET NIMCET 2017 PYQ

Solution

Mean Deviation from mean is Zero.

MCA NIMCET PYQ
Let $X$ and $Y$ be two independent identically distributed Bernoulli random variables with common probability mass function: $P(X=1)=\frac{1}{2}$ and $P(X=0)=\frac{1}{2}$. If $Z=XY$, then the distribution of $Z$ is:





Go to Discussion

MCA NIMCET Previous Year PYQ MCA NIMCET NIMCET 2026 PYQ

Solution

Given,

$Z=XY$

Now, $Z=1$ only when both $X=1$ and $Y=1$.

Since $X$ and $Y$ are independent,

$P(Z=1)=P(X=1,Y=1)$

$P(Z=1)=P(X=1)P(Y=1)$

$P(Z=1)=\frac{1}{2}\times \frac{1}{2}$

$P(Z=1)=\frac{1}{4}$

Now,

$P(Z=0)=1-P(Z=1)$

$P(Z=0)=1-\frac{1}{4}$

$P(Z=0)=\frac{3}{4}$

Therefore, $Z$ follows Bernoulli distribution with

$P(Z=1)=\frac{1}{4}$ and $P(Z=0)=\frac{3}{4}$

Correct answer is option $4$.



MCA NIMCET PYQ
For a given sample, the computed values of the variance and fourth central moment are $3$ and $63$ respectively. Then the underlying frequency distribution is classified as:





Go to Discussion

MCA NIMCET Previous Year PYQ MCA NIMCET NIMCET 2026 PYQ

Solution

Given,

Variance $=3$

So,

$\mu_2=3$

Fourth central moment is:

$\mu_4=63$

Coefficient of kurtosis is:

$\beta_2=\frac{\mu_4}{\mu_2^2}$

Substitute the values:

$\beta_2=\frac{63}{3^2}$

$\beta_2=\frac{63}{9}$

$\beta_2=7$

For a normal or mesokurtic distribution,

$\beta_2=3$

Here,

$\beta_2=7>3$

So, the distribution is leptokurtic.Given,

$Z=XY$

Now, $Z=1$ only when both $X=1$ and $Y=1$.

Since $X$ and $Y$ are independent,

$P(Z=1)=P(X=1,Y=1)$

$P(Z=1)=P(X=1)P(Y=1)$

$P(Z=1)=\frac{1}{2}\times \frac{1}{2}$

$P(Z=1)=\frac{1}{4}$

Now,

$P(Z=0)=1-P(Z=1)$

$P(Z=0)=1-\frac{1}{4}$

$P(Z=0)=\frac{3}{4}$

Therefore, $Z$ follows Bernoulli distribution with

$P(Z=1)=\frac{1}{4}$ and $P(Z=0)=\frac{3}{4}$

Correct answer is option $4$.




MCA NIMCET


Online Test Series,
Information About Examination,
Syllabus, Notification
and More.

Click Here to
View More

MCA NIMCET


Online Test Series,
Information About Examination,
Syllabus, Notification
and More.

Click Here to
View More

Limited Seats
× Aspire MCA Promotion

Game Changer NIMCET Test Series 2026

Boost your preparation with mock tests, analysis and rank-focused practice.

JOIN NOW
Ask Your Question or Put Your Review.

loading...